OpenStudy (anonymous):

What am I doing wrong? DIFF EQ Initial Value Problem: \[\frac{ dy }{ dx } + \sqrt{1+t^2} e ^{-1}y = 0 ; y(0)=1\] First, I found the Integrating Factor: \[e^{\int\limits_{}^{}\sqrt{1+t^2}e ^{-1}dt}\] = \[e ^{{(\sqrt{1+t^2}+\sinh ^{-1}(t)})/2e}\]

OpenStudy (anonymous):

And I thought that I should have \[y(t) = ce ^{{(\sqrt{1+t^2}+\sinh ^{-1}(t)})/2e}\] but my book disagrees... any suggestions?

OpenStudy (phi):

wolfram came up with a different answer for the integral http://www.wolframalpha.com/input/?i=int+sqrt%281%2Bx%5E2%29%2Fe+dx

OpenStudy (anonymous):

I keep getting disconnected while trying to type responses.

OpenStudy (anonymous):

The book gives the answer \[y(t) = \exp[\int\limits_{0}^{t}\sqrt{1+s^2}e ^{-1}ds]\] my questions are, why are the bounds 0 to t. Is that because the intial value is y(0)=1? So if it had been y(3)=1 for example, the bounds would be 3 to t?

OpenStudy (anonymous):

and also, why is it in terms of s, when the problem started with y's and t's?

OpenStudy (anonymous):

None of these are explained in this book, so my guess is they are assumptions on something I should remember from calculus.

OpenStudy (phi):

the s is a dummy value. I think they though the integral was to hard to solve, so they leave it in integral form

OpenStudy (anonymous):

ya thats what I figure, it was a beast. What about the bounds on it though?

OpenStudy (phi):

I would have to think about the bounds.

OpenStudy (anonymous):

Dont worry about it, I have to head to class in a little while, for a different math class haha. Abstract Algebra is much more to my liking. Thanks for the assistance as always though!